Write each function in terms of its co function. Assume all angles in which an unknown appears are acute angles.
cos(t-5)=

Answer by stanbon(75887)

Write each function in terms of its co function. Assume all angles in which an unknown appears are acute angles.
cos(t-5)= sin(90-(t-5)) = sin(95-t)
======================================
Cheers,
Stan H.
===============================